In 1932 , I was as old as the number formed by the last two digits of my birth year. When I mentioned the interesting coincidence to my grandfather, he surprised me by saying that the same applied to him also. If so find my year of birth and also my grandfather.
1
Expert's answer
2019-05-10T11:47:40-0400

It is reasonable to assume that my grandfather was born in the nineteenth century (he was a devout catholic) since I was alive in 1932. Therefore, his age should be a solution of the equation

1932-(1800+y)=y (0<=y<100<=132)

1932-1800-y=y

1932-1800=2y

132=2y

y=66


Because this equation has only one solution, I must be born in the twentieth century. Thus, my age should be a solution of the equation

1932-(1900+x)=x (0<=x<=32<100)

1932-1900-x=x

1932-1900=2x

32=2x

x=16


From these findings, we conclude that my year of birth is 1916, and my grandfather’s year of birth is 1866.